P12.037 The stresses shown act at a point on the free surface of a stressed body. Normal and shear stress magnitudes acting oAnswers: On= MPa o MPа Int = MPa
Answer #1

- бу омро бz - 98 zy - Say=12 6 = 98 тра, бу = - 20 тра, аяу = 12 оро В : 52°, 9 - 90 - 51 - 33° ) normal stress (60) бо - бя3 67 = 62-6y sin 20 - cay Cos20 - 98+20 - Sin 66 - 72 cos 66 67 = 53.899 -29.285 - 24.61 mpa 3) maximum shear stress (Ant) to